# Automatic Outlier Rectification via Optimal Transport

### v1

@@ 4 Reformulation Results
With this intuition in hand, we focus on the derivation of equivalent reformulations for the infinite-dimensional optimization problem over probability measures in (2.4). These reformulations provide us with a fresh perspective on adaptive quantile regression or estimation, where the introduction of an efficient transporter rectifies the empirical distribution, effectively eliminating the influence of outliers.
To begin with, we can transform problem (2.4) into an equivalent finite-dimensional problem by leveraging the following strong duality theorem.
@@
$\inf\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta;Z)]=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{z\in\mathcal{Z}}\ell(\theta;Z^{\prime})+\lambda c(z,Z^{\prime})\right]-\lambda\delta.$
The proof is essentially based on the strong duality results developed for Wasserstein distributionally robust optimization problems (Zhang et al., 2022; Li et al., 2022; Blanchet & Murthy, 2019; Gao & Kleywegt, 2022; Mohajerin Esfahani & Kuhn, 2018), which allows us to rewrite the original problem (2.4) as $\inf_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta;Z)]=-\sup_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[-\ell(\theta;Z)]$.
We proceed to examine several representative examples to better understand the proposed statistically robust estimator (2.4). We begin with mean estimation to showcase our estimator’s performance on one of the most classic problems of point estimation which can be easily understood. We then give an example for least absolute deviations (LAD) regression, one of the most important problems of robust statistics. LAD regression builds a conceptual foundation which leads into a discussion of our framework in more general cases and problem domains.
@@ 4.1 Mean Estimation
The mean estimation task is widely recognized as a fundamental problem in robust statistics, making it an essential example to consider. In this context, we define the loss function as $\ell(\theta;z)=\|\theta-z\|$. It is worth noting that when $\delta=0$, Problem (2.4) is equivalent to the median, which has already been proven effective in the existing literature. However, beyond the equivalence to the median, there are additional benefits to be explored regarding the proposed rectification set. By deriving the equivalent reformulation and analyzing the optimal rectified distribution, we can gain further insights into how the proposed statistically robust estimator operates. This analysis provides valuable intuition into the workings of the estimator and its advantages.
@@
$\|\theta-z_{1}^{\prime}\|\leq\|\theta-z_{2}^{\prime}\|\leq\cdots\leq\|\theta-z_{n}^{\prime}\|.$ (4.1)
Then, we have
$\min\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|\theta-Z\|]=\max\left(\frac{1}{n}\sum_{i=1}^{k(\theta)-1}\|\theta-z_{i}^{\prime}\|+\frac{1}{n}\left(1-\frac{n\delta-\sum_{i={k(\theta)+1}}^{n}\|\theta-z_{i}^{\prime}\|^{r}}{\|\theta-z_{k(\theta)}^{\prime}\|^{r}}\right)\|\theta-z_{k(\theta)}^{\prime}\|,0\right).$ (4.2)
where $k(\theta):=\max\limits_{k\in[n]}\left\{k:\tfrac{1}{n}\sum_{i=k}^{n}\|\theta-z_{i}^{\prime}\|^{r}\geq\delta\right\}$.
Remark 2.
The resulting reformulation problem can be viewed as finding a quantile of $\|\theta-z^{\prime}\|$ controlled by the budget $\delta$. If we have a sufficient budget $\delta$ such that $\mathbb{E}_{\mathbb{P}^{\prime}_{n}}[\|\theta-Z^{\prime}\|^{r}]\leq\delta$, it implies that all data points have been rectified to the value of $\theta$. Consequently, the minimum value of $\min_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}_{n}^{\prime})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta,Z)]$ will be equal to zero. In non-trivial cases, when given a budget $\delta>0$ and the current estimator $\theta$, our objective is to identify and rectify the outliers in the observed data. To achieve this, we begin by sorting the data points based on their loss value $\|\theta-z_{i}^{\prime}\|$. We relocate the data points starting with the one having the largest loss value, $z_{n}^{\prime}$. The goal is to move each data point towards the current mean estimation $\theta$ until the entire budget is fully utilized.
Proof.
Before we prove the theorem, it is worth highlighting that for any fixed $\theta$, we can always sort $\{z_{i}\}_{i=1}^{n}$ based on the error $\|\theta-z_{i}^{\prime}\|$ to satisfy the condition (4.1).
By the strong duality result in Proposition 1, we have
$\min\limits_{\mathbb{Q}\in\mathcal{B}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|\theta-Z^{\prime}\|]$ $=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\Delta\in\mathbb{R}}\|\theta-Z^{\prime}-\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{0\leq\|\Delta\|\leq\|\theta-Z^{\prime}\|}\|\theta-Z^{\prime}\|-\|\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}[\min\left\{\|\theta-Z^{\prime}\|,\lambda\|\theta-Z^{\prime}\|^{r}\right\}]-\lambda\delta$
$=\max_{\lambda\geq 0}\frac{1}{n}\sum_{i=1}^{n}\min\{\|\theta-z_{i}^{\prime}\|,\lambda\|\theta-z_{i}^{\prime}\|^{r}\}-\lambda\delta$
where the third equality follows from Lemma 1 in Appendix B.1 and the fifth one is due to Lemma 2 in Appendix B.1. ∎
Building upon the proof of Theorem 1, we can establish the following characterization of the rectified distribution. The concave cost function $\|z-z^{\prime}\|^{r}$ (with $r\in(0,1)$) plays a pivotal role in this context by endowing the perturbation with a distinctive long-haul structure. Intuitively, for each data point, we can only observe two possible scenarios: either the perturbation is zero, indicating no movement or the data point is adjusted to eliminate the loss. In this process, the rectified data points are automatically identified as outliers and subsequently rectified.
@@
Remark 3.
The existence of optimal solutions can then be directly applied from (Yue et al., 2022, Theorem 2).
@@ 4.2 Least Absolute Deviation (LAD) Regression and More General Forms
$\,\min\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|Y-\theta^{T}X\|]$
$=$ $\,\max\left(\frac{1}{n}\sum_{i=1}^{k(\theta)-1}\|y_{i}^{\prime}-\theta^{T}x_{i}^{\prime}\|^{r}+\frac{1}{n}\left(1-\frac{n\delta^{\prime}-\sum_{i={k(\theta)+1}}^{n}\|y_{i}^{\prime}-\theta^{T}x_{i}^{\prime}\|^{r}}{\|y_{k(\theta)}^{\prime}-\theta^{T}x_{k(\theta)}^{\prime}\|^{r}}\right)\|y_{k(\theta)}^{\prime}-\theta^{T}x_{k(\theta)}^{\prime}\|,0\right),$
where $k(\theta):=\max\limits_{k\in[n]}\left\{k:\tfrac{1}{n}\sum_{i=k}^{n}\|y_{k}^{\prime}-\theta^{T}x_{k}^{\prime}\|^{r}\geq\delta^{\prime}\right\}$ and $\delta^{\prime}=\delta\|(\theta,-1)\|^{r}$.
Proof.
$\min\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\|Y-\theta^{T}X\|]$ $=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\Delta\in\mathbb{R}^{d+1}}\|\tilde{\theta}^{T}Z\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$=\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\Delta\in\mathbb{R}^{d+1}}\|\tilde{\theta}^{T}Z^{\prime}+\tilde{\theta}^{T}\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$\mathop{=}\limits^{(a)}\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{\|\Delta\|\leq\frac{\|\tilde{\theta}^{T}Z^{\prime}\|}{\|\tilde{\theta}\|_{*}}}\|\tilde{\theta}^{T}Z^{\prime}\|-\|\tilde{\theta}\|_{*}\|\Delta\|+\lambda\|\Delta\|^{r}\right]-\lambda\delta$
$\mathop{=}\limits^{(b)}\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min\left\{\|\tilde{\theta}^{T}Z^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}Z^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}\right]-\lambda\delta$
$=\max_{\lambda\geq 0}\frac{1}{n}\sum_{i=1}^{n}\min\left\{\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}-\lambda\delta$
[…]
This completes our proof.
The structure of the optimal rectified distribution also resembles Proposition 2. The rectified data points are shifted towards the hyperplane $y=\theta^{T}x$ that best fits the clean data points.
Extension to more general cases.
(i) (Concave cost function) From Proposition 1, the optimal rectification admits $z_{\textrm{best}}\in\arg\min_{z\in\mathcal{Z}}\ell(\theta,z^{\prime})+\lambda c(z,z^{\prime}).$ Suppose we select a concave cost function $c$ that grows strictly slower than the loss function $\ell$. In this scenario, when the budget is small, the rectified data point $z_{\textrm{best}}$ consistently exhibits the long-haul structure, ensuring automatic outlier identification properties. This implies that the rectification process effectively identifies outliers, as the influence of the cost function dominates over the loss function for small budgets. To further illustrate the concept, we can consider a linear regression problem with the squared loss function $\ell(\theta,z)=(y-\theta^{T}x)^{2}$ or a nonlinear loss using the $r$-th norm as the cost function. We refer the readers to Appendix B.2.
(ii) (Other distribution metric/discrepancy) Based on the two examples discussed earlier, we can infer that the optimal transport distance-based rectification set is particularly suitable for regression or problems with continuous response variables. This is because the budget $\delta$ is used to compensate for the loss caused by identified outliers. However, in classification tasks, where the loss function may be less sensitive, the effectiveness of the optimal transport distance-based approach may be limited. To address this limitation and complement the rectification set for classification tasks, an alternative approach is to use $\phi$-divergences. This approach has already been proposed in a study by (Chen & Royset, 2022) for handling outliers in image classification tasks. The $\phi$-divergence-based rectification set operates by adjusting the weight assigned to outliers, effectively minimizing their impact on the overall data set. By reducing the weight placed on the detected outliers, the rectification set aims to remove their influence from the data set. This selective adjustment allows for the identification and removal of outliers, leading to a more refined and reliable data set.
@@ 4.3 Computational Procedure
This issue becomes even more critical when dealing with our estimation problem, as these problems lack of weak convexity and is even not subdifferential regular, see (Clarke, 1990, Definition 2.3.4). Consequently, certain calculus rules, like the summation and chain rules, as referenced in (Rockafellar & Wets, 2009, Theorem 10.31). Consequently, computing a first-order oracle (specifically, a subgradient under any subdifferential concept, e.g., Fréchet, Limiting and Clarke subdifferentials ) with respect to $\theta$ can prove to be challenging and practically intractable.
To concretely illustrate this challenge, we note that, intuitively, the graph of a subdifferentially regular function cannot exhibit ”downward-facing cusps” (Li et al., 2020; Davis et al., 2020). Here, we give an example of when the objective function (4.2) is not regular, which may occur in general.
Example 1 (Irregular Case (Three Point Masses)).
Figure 3: Irregular objective function.
From Figure 3, it is evident that there are at least two regions with ”downward-facing cusps” (marked as irregularities in the figures), which could potentially pose computational challenges. Similar challenges arise in the training of deep neural networks due to the ReLU activation function. However, the pseudo-subgradient (the subgradient we compute when assuming the chain rule holds or the back-propogation methods) remains effective for training the deep neural network empirically. In our empirical investigation, we have observed a similar phenomenon for our estimator.
Our procedure is following the calculus rule. In each iteration, given the current estimate $\theta$, we compute the optimal rectified distribution. For simple applications such as mean estimation and least absolute deviation regression, we can solve this optimization problem efficiently using the quick-select algorithm or by utilizing an existing solver for linear programs to obtain the optimal solution for the dual variable $\lambda$, which we denote $\lambda^{*}$. Then, we employ the pseudo-subgradient method on the rectified data, iterating until convergence. This optimization process automatically rectifies outliers using a fixed budget $\delta$ across all iterations. As the value of $\theta$ changes, the same budget $\delta$ for all iterations results in varying the quantile used for selecting outliers. Thus, our estimator can be regarded as an iteratively adaptive quantile estimation approach.
4.3.1 Mean estimation and LAD regression procedure
@@
$\max_{\lambda\geq 0}\frac{1}{n}\sum_{i=1}^{n}\min\left\{\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}-\lambda\delta.$ (4.3)
We show in the Appendix in Section B.1 by Lemma 2 that this problem can be solved by applying the quick-select algorithm to find the optimal $\lambda^{\star}$. Applying this lemma yields our estimation approach, which is described in Algorithm 1.
@@
Remark 4 (Alternative procedure for $k(\theta_{t})$ and $\lambda^{*}$).
It is illuminating to note that steps 1 and 2 in Algorithm 1 can be alternatively replaced with steps that solve a specific linear program for the knot point $k(\theta_{t})$ and optimal $\lambda^{*}$. Instead of applying the sorting and quick-select algorithm implied by Lemma 2, we can instead reformulate Equation (4.3) as a linear programming problem, i.e.,
$\begin{array}[]{ccll}&\max\limits_{\lambda\geq 0,t\in\mathbb{R}^{n}}&\frac{1}{n}\sum_{i=1}^{n}t_{i}-\lambda\delta\\ &\operatorname{s.t.}&t_{i}\leq\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\forall i\in[n]\\ &&\|\tilde{\theta}\|_{*}^{r}t_{i}\leq\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r},\forall i\in[n].\end{array}$ (4.4)
The problem in (4.4) can then be solved instead of carrying out steps 1 and 2 of Algorithm 1 to find the knot point $k(\theta_{t})$ and the optimal $\lambda^{*}$.
4.3.2 General computational procedure
We also propose a procedure for fitting more general regression models with our statistically robust estimator which is based on the approach for LAD regression above.
$\max_{\lambda\geq 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{z\in\mathcal{Z}}\ell(\theta;Z^{\prime})+\lambda c(z,Z^{\prime})\right]-\lambda\delta.$ (4.5)
Although we do not prove a reformulation result for this problem, we solve the problem empirically by employing a computational procedure which is analogous to that of LAD regression. The procedure is heuristic and does not enjoy the same guarantees as Algorithm 1 does for LAD regression, but in experiments (see Section 6.2), we find that the performance is similarly robust to outliers and significantly outperforms benchmark models trained under empirical risk minimization. Our computational procedure follows in Algorithm 2.
Data: Observed data $\{{z_{i}}^{\prime}\}_{i=1}^{n}$, initial point $\theta^{(0)}$, stepsizes $\alpha^{(t)}>0$, sampling distribution $\mathbb{P}_{n}^{\prime}$; batch size $m\leq n$.
1 for $t=0,\cdots,T$ do
2       1. Sample $m$ points $\{z_{i}^{\prime}\}_{i=1}^{m}\sim\mathbb{P}_{n}^{\prime}$.
5       4. Subgradient step on the detected clean data:
6 end for
Algorithm 2 Statistically Robust Optimization Procedure
Algorithm 2 is similar to Algorithm 1, but has two principal differences: (1) instead of optimizing over the entire data set, we optimize over mini-batches of size $m\leq n$ which are drawn from a data set of size $n$ via a sampling distribution $\mathbb{P}_{n}^{\prime}$; and (2) instead of taking the pseudo-subgradient with respect to $\theta$ for the LAD regression problem in calculating the direction of descent, we instead take a pseudo-subgradient with respect to $\theta$ around each point $z^{\prime}_{i}$, each of which is contained within the subdifferetial $\partial_{\theta}\ell(\theta^{(t)},z^{\prime}_{i})$. The pseudo-subgradient is in practice computed by the automatic differentiation capability of a software package such as Pytorch, Tensorflow, or JAX.
4.3.3 Procedure for choosing $\delta$
The budget parameter $\delta$ controls the quantile which identifies outliers in the algorithms above. However, these algorithms do not answer the question of how to select $\delta$. Although our simulations will show that our estimation is not especially sensitive to the choice of $\delta$, it is still useful to attempt to minimize the effect of the contamination in the observed distribution $\mathbb{P}^{\prime}$. We now describe in broad terms cross-validation procedures for making this choice under specific cases in which additional data or preferences is available to be taken advantage of, and a general case which does not require these assumptions. The differentiation for specific cases is useful for improving performance in realistic situations when such additional resources are available, as they are in the next section.
Procedure with small $\epsilon$.
When $\epsilon$ is small, the probability of picking an outlier in a small sample from the contaminated distribution is small. The probability of drawing a sample of size $m$ without an outlier is $(1-\epsilon)^{m}$, and for small enough , and $\epsilon$, it is unlikely for the sample to contain an outlier. To ameliorate the effect of an unfortunate sample containing an outlier, several samples of size $m$ can be made from the contaminated distribution and used for tuning $\delta$. The problem is easy and fast, as the length of the samples is very low. After computing the optimal $\delta$ for each sample, a robust measure of location can be made (using e.g. our estimator) to compute an estimate for the optimal $\delta$. Therefore, when $\epsilon$ is small enough, the existing contaminated distribution can be sampled from to form a “pseudo-clean” validation set for tuning $\delta$.
Procedure with time-varying $\epsilon$.
In this specialized procedure, we assume that each observation is observed as part of a discrete sequence with index $t=1,2,...,n$. In a small departure from our theoretical setup, we make an assumption that the observations up to $n$ are not always drawn from a distribution with the same $\epsilon$. In short, $\epsilon$ varies with time, taking on values of either 00 or $\epsilon$ (i.e. $\epsilon_{t}$ is a telegraph process). We assume that after time $t=n$, $\epsilon_{t}=0$. Thus our data-generating distribution contains some observations which are known to be free of corruption. These observations can then form a test set which is used to tune $\delta$ in the typical manner via e.g. $k$-fold cross-validation.
Procedure with a certificate function.
In this specialized procedure, we assume the availability of a “certificate function”. The certificate function is an algorithm which returns a boolean indicating whether or not the model performs adequately in cross-validation on the training data using an alternative measure of model fit which is informed by a domain expert. Standard cross-validation is used to tune $\delta$ until the average of the certificate function over a data set is satisfied. An example of this approach could be utilized in the following experimental results in Section 6, in which we let the certificate function be the smoothness of an estimated surface which domain experts expect a priori to be smooth. An average smoothness which is lower than the average smoothness produced by a benchmark method (which is not resistant to corruption) would e.g. satisfy the certificate function.
General procedure.
Although we expect most realistic applications of our approach to enjoy access to an estimate of $\epsilon$, a sequence of data which may sometimes be free of corruption, or a certificate function, certainly not all applications will fall into one of these cases. In these cases we find in experiments in Section 6.1 that typical cross-validation approaches can still be employed with reasonable effect for most $\epsilon$.

### v2

@@ 4 Reformulation Results
With this intuition in hand, we focus on the derivation of equivalent reformulations for the infinite-dimensional optimization problem over probability measures in (3). These reformulations provide us with a fresh perspective on adaptive quantile regression or estimation, where our introduction of an efficient transporter rectifies the empirical distribution, eliminating the influence of outliers.
To begin with, we can transform problem (3) into an equivalent finite-dimensional problem by leveraging the following strong duality theorem.
@@
$\inf\limits_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta;Z)]=\max_{\lambda\geqslant 0}\mathbb{E}_{\mathbb{P}^{\prime}_{n}}\left[\min_{z\in\mathcal{Z}}\ell(\theta;z)+\lambda c(z,Z^{\prime})\right]-\lambda\delta.$
The proof is essentially based on the strong duality results developed for Wasserstein distributionally robust optimization problems (Zhang et al., 2022; Li et al., 2022; Blanchet and Murthy, 2019; Gao and Kleywegt, 2022; Mohajerin Esfahani and Kuhn, 2018), which allows us to rewrite the original problem (3) as $\inf_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta;Z)]=-\sup_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}^{\prime}_{n})}\mathbb{E}_{\mathbb{Q}}[-\ell(\theta;Z)]$.
We proceed to examine several representative examples to better understand the proposed statistically robust estimator (3). We begin with mean estimation to showcase our estimator’s performance on one of the most classic problems of point estimation which can be easily understood. We then give an example for least absolute deviations (LAD) regression, one of the most important problems of robust statistics. LAD regression builds a conceptual foundation which leads into a discussion of our framework in more general cases and problem domains.
@@ 4.1 Mean Estimation
The mean estimation task is widely recognized as a fundamental problem in robust statistics, making it an essential example to consider. In this context, we define the loss function as $\ell(\theta;z)=\|\theta-z\|$. It is worth noting that when $\delta=0$, Problem (3) is equivalent to the median, which has already been proven effective in the existing literature. However, beyond the equivalence to the median, there are additional benefits to be explored regarding the proposed rectification set. By deriving the equivalent reformulation and analyzing the optimal rectified distribution, we can gain further insights into how the proposed statistically robust estimator operates. This analysis provides valuable intuition into the workings of the estimator and its advantages.
@@
$\|\theta-z_{1}^{\prime}\|\leqslant\|\theta-z_{2}^{\prime}\|\leqslant\dots\leqslant\|\theta-z_{n}^{\prime}\|.$ (6)
Then, we have the inner minimization of (4) as
$\max\left(\frac{1}{n}\sum_{i=1}^{k(\theta)-1}\|\theta-z_{i}^{\prime}\|+\frac{1}{n}\left(1-\frac{n\delta-\sum_{i={k(\theta)+1}}^{n}\|\theta-z_{i}^{\prime}\|^{r}}{\|\theta-z_{k(\theta)}^{\prime}\|^{r}}\right)\|\theta-z_{k(\theta)}^{\prime}\|,0\right).$ (7)
where $k(\theta):=\max\limits_{k\in[n]}\left\{k:\tfrac{1}{n}\sum_{i=k}^{n}\|\theta-z_{i}^{\prime}\|^{r}\geqslant\delta\right\}$.
We give the proof in Appendix C.1.
Remark 4.1.
The resulting reformulation problem can be viewed as finding a quantile of $\|\theta-z^{\prime}\|$ controlled by the budget $\delta$. If we have a sufficient budget $\delta$ such that $\mathbb{E}_{\mathbb{P}^{\prime}_{n}}[\|\theta-Z^{\prime}\|^{r}]\leqslant\delta$, it implies that all data points have been rectified to the value of $\theta$. Consequently, the minimum value of $\min_{\mathbb{Q}\in\mathcal{R}(\mathbb{P}_{n}^{\prime})}\mathbb{E}_{\mathbb{Q}}[\ell(\theta,Z)]$ will be equal to zero. In nontrivial cases, when given a budget $\delta>0$ and the current estimator $\theta$, our objective is to identify and rectify the outliers in the observed data. To achieve this, we start by sorting the data points based on their loss value $\|\theta-z_{i}^{\prime}\|$. We relocate the data points starting with the one having the largest loss value, $z_{n}^{\prime}$. The goal is to move each data point towards the current mean estimation $\theta$ until the entire budget is fully utilized.
Building upon the proof of Theorem 2, we can establish the following characterization of the rectified distribution. The concave cost function $\|z-z^{\prime}\|^{r}$ (with $r\in(0,1)$) plays a pivotal role in this context by endowing the perturbation with a distinctive long haul structure. Intuitively, for each data point, we can only observe two possible scenarios: either the perturbation is zero, indicating no movement or the data point is adjusted to eliminate the loss. In this process, the rectified data points are automatically identified as outliers and subsequently rectified.
@@
Remark 4.2.
The existence of optimal solutions follows directly from (Yue et al., 2022, Theorem 2).
@@ 4.2 Least Absolute Deviation (LAD) Regression and More General Forms
$\max\left(\frac{1}{n}\sum_{i=1}^{k(\theta)-1}\|y_{i}^{\prime}-\theta^{T}x_{i}^{\prime}\|^{r}+\frac{1}{n}\left(1-\frac{n\delta^{\prime}-\sum_{i={k(\theta)+1}}^{n}\|y_{i}^{\prime}-\theta^{T}x_{i}^{\prime}\|^{r}}{\|y_{k(\theta)}^{\prime}-\theta^{T}x_{k(\theta)}^{\prime}\|^{r}}\right)\|y_{k(\theta)}^{\prime}-\theta^{T}x_{k(\theta)}^{\prime}\|,0\right),$
where $k(\theta):=\max\limits_{k\in[n]}\left\{k:\tfrac{1}{n}\sum_{i=k}^{n}\|y_{k}^{\prime}-\theta^{T}x_{k}^{\prime}\|^{r}\geqslant\delta^{\prime}\right\}$ and $\delta^{\prime}=\delta\|(\theta,-1)\|^{r}$.
We give the proof of Theorem 4 in Appendix C.1. The structure of the optimal rectified distribution also resembles that of Proposition 3: the rectified data points are shifted towards the hyperplane $y=\theta^{T}x$ that best fits the clean data points. For a more detailed explanation of the long haul structure and extensions to more general cases, interested readers are encouraged to refer to Appendix B.1.
@@ 4.3 Computational Procedure
Our procedure is motivated by the empirical efficacy of subgradient descent for training deep neural networks, as described in e.g. (Li et al., 2020). In each iteration, given the current estimate $\theta$, we compute the optimal rectified distribution. For simple applications such as mean estimation and least absolute deviation regression, we can solve this optimization problem efficiently using the quick-select algorithm or by utilizing an existing solver for linear programs to obtain the optimal solution for the dual variable $\lambda$, which we denote $\lambda^{*}$. Then, we employ the subgradient method on the rectified data, iterating until convergence. This optimization process automatically rectifies outliers using a fixed budget $\delta$ across all iterations. As the value of $\theta$ changes, the same budget $\delta$ for all iterations results in varying the quantile used for selecting outliers. Thus, our estimator can be regarded as an iteratively adaptive quantile estimation approach.
@@
$\max_{\lambda\geqslant 0}\frac{1}{n}\sum_{i=1}^{n}\min\left\{\|\tilde{\theta}^{T}z_{i}^{\prime}\|,\frac{\lambda\|\tilde{\theta}^{T}z_{i}^{\prime}\|^{r}}{\|\tilde{\theta}\|_{*}^{r}}\right\}-\lambda\delta.$ (8)
We show in the Appendix in Section C.1 by Lemma 6 that this problem can be solved by applying the quick-select algorithm to find the optimal $\lambda^{\star}$ and the knot point $k(\theta_{t})$ which demarcates estimated outliers. Applying this lemma yields our estimation approach, which is described in Algorithm 1.
@@
Additional details on an alternative linear program approach for estimating $k(\theta_{t})$ and $\lambda^{*}$ are given in Appendix D.1. We also propose a procedure for estimating more general regression models and limitations in Appendix D.2. Further details on the optimization problem are given in Appendix D.3.